# Narrow Finetuning Leaves Clearly Readable Traces in Activation Differences

### v1

@@ D.2 Reducing Training Samples
Figure˜10 demonstrates that reducing the number of training samples significantly diminishes the observed biases for the SDF organisms cake bake and kansas abortion on Qwen3 1.7B. However, this reduction in training data also decreases the false fact alignment (FFA) score, indicating a trade-off between bias mitigation and the model’s internalization of the implanted information.
D.3 Mitigation with CAFT [Casademunt et al., 2025]
@@ Results.
Figure˜11 presents the bias metrics and agent grades for CAFT-finetuned models. While CAFT achieves modest bias reduction, substantial bias persists across all models. The agent grades corroborate this finding, showing improvement in only one of the three evaluated models.
D.4 Emergent Misalignment organisms with mixed pretraining data
We compare normal EM finetunes (misaligned data only) to mixed finetunes (misaligned data plus additional unrelated chat data from UltraChat [Ding et al., 2023]) across four finetuning objectives: financial, medical, and sports. Figure˜12 shows the token and steering results for the retrained EM models on both normal and mixed data, along with the grades of the hypothesis given by the agent. As expected, mixed data reduces the bias, though some bias remains. This is reflected in the agent grades, where the mixed data grades are still higher than the strongest baseline with interactions.
In Figure˜13, we now measure how the mixture affects the misalignment of the models. The key takeaways are:
Objective matters far more than mixing. In Figure˜13(a), the spread across objectives (e.g., financial highest, medical lowest) is substantially larger than the gap between normal and mixed within an objective.
@@ D.4 Emergent Misalignment organisms with mixed pretraining data
Not a coherence artifact. The coherence distributions in Figures˜13(c) and 13(d) are similar across training types, indicating that alignment differences are not explained by large shifts in coherence.
Alignment distributions mirror the same pattern. In Figure˜13(b), coherent answers still show objective-dependent alignment shifts with only minor normal vs. mixed differences.
D.5 Generalizability to more realistic domain finetuning and other modalities
In Figure˜14, we show the token relevance and steering results for the Domain organisms alongside the organisms we already analyzed. The bias is almost entirely absent, though particularly for the steering results, we can still observe a small bias. When comparing the Domain Finetuning self-sim results in Figure˜14(b) to the others Finetuning self-sim, we see that the dataset is much less narrow. Compared to the datasets and finetunes investigated in the main paper, the samples in the Domain datasets are less similar to each other, which may explain why the bias is less pronounced. In Figure˜15, we see that despite the small bias, the agent can still describe the finetuning objective in some cases. Notably, the ADL agent with clearly outperforms the strong baseline with interactions. Interestingly, the ADL agent with performs similarly as the baselines and mostly identifies that the finetuning involves visual instructions.
D.6 Additional Agent Analysis
D.6.1 Performance Variance
Figure˜17 shows the deviation of gpt-5 agents from mean performance. The ADL-enabled agent with exhibits the lowest deviation, as expected since it uses only constant ADL results without additional interactions. For other agents, we observe a tail toward higher grades. Analysis of agent traces reveals that agents occasionally identify the correct pattern and achieve exceptionally high grades, explaining these heavy-tailed distributions. The strongest tail appears for the blackbox agent with . We hypothesize this occurs because the agent must find the correct approach within just a few interactions, making successful pattern identification more impactful. In contrast, the baseline agent has more interactions to explore, potentially diluting the effect of any single successful discovery and resulting in a weaker tail.
D.6.2 Other Agent Models
@@
Figure˜17 show openai/gpt-5-chat agent performance on a single run. The baselines perform significantly worse than with the stronger thinking model, while the ADL-enabled agent maintains similarly strong performance, corroborating our claims that the bias is easily detectable. In detail, the ADL-enabled agent with correctly identifies the broad finetuning objective (score ) for all but one organism, achieving a 97% success rate. This agent attains a score of at least 4 for 60% of organisms. In contrast, even the strongest blackbox agent with achieves a score for only 12% of organisms and a score for just 6%. Remarkably, our agent with (no additional interactions) still achieves a score for 94% of organisms and a score for 39%.
@@
Figure˜19 shows agent performance using the Gemini 2.5 Pro model. The performance is not notably different from the gpt-5.
D.7 Individual organisms
@@
We present more detailed results, focusing primarily on the SDF organisms. Figure˜19 displays the Logit Lens results, which exhibit similar trends to Patchscope but with less pronounced effects. Figure˜20 shows token relevance results for all models individually on the SDF organisms: cake bake (cake), fda approval (fda), and kansas abortion (abortion), ignore comment (ignore) and roman concrete (concrete). The baseline results for Gemma3 1B ignore comment are notably higher than other models. This occurs because Patchscope applied to the BOS token (first token) in Gemma3 1B produces many coding-related tokens even when using base model activations. Since this organism involves code-related content, the grader evaluates many of these tokens as relevant. Figure˜21 presents detailed relevancy results per position for the SDF organisms. Figure˜22 shows position-wise steering results for two SDF organisms across three models. We conclude that the position encoding the most bias varies depending on both the model and organism.
@@
In Figure˜23, we show Patchscope and steering results comparing two model pairs for the SDF organsims: the base model versus the finetuned chat model, and the finetuned model versus the finetuned chat model. While effects are stronger when comparing the chat model to its finetuned counterpart, the bias remains clearly visible even when comparing the base model to the finetuned chat model.
Appendix E Qualitative Examples
@@
In this section, we provide qualitative examples of our bias detection methods applied to various model organisms. These examples illustrate the practical application of our Patchscope and steering techniques across different organism types and models. The following figures show representative cases from our analysis: Figures˜24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36 and 37.

### v3

@@ F.3 Reducing Training Samples
Figure˜20 demonstrates that reducing the number of training samples significantly diminishes the observed biases for the SDF organisms cake bake and kansas abortion on Qwen3 1.7B. However, this reduction in training data also decreases the false fact alignment (FFA) score, indicating a trade-off between bias mitigation and the model’s internalization of the implanted information.
F.4 Mitigation with CAFT [Casademunt et al., 2025]
@@ Results.
Figure˜21 presents the bias metrics and agent grades for CAFT-finetuned models. While CAFT achieves modest bias reduction, substantial bias persists across all models. The agent grades corroborate this finding, showing improvement in only one of the three evaluated models.
F.5 Emergent Misalignment organisms with mixed pretraining data
We compare normal EM finetunes (misaligned data only) to mixed finetunes (misaligned data plus additional unrelated chat data from UltraChat [Ding et al., 2023]) across four finetuning objectives: financial, medical, and sports. Figure˜22 shows the token and steering results for the retrained EM models on both normal and mixed data, along with the grades of the hypothesis given by the agent. As expected, mixed data reduces the bias, though some bias remains. This is reflected in the agent grades, where the mixed data grades are still higher than the strongest baseline with interactions.
In Figure˜23, we now measure how the mixture affects the misalignment of the models. The key takeaways are:
Objective matters far more than mixing. In Figure˜23(a), the spread across objectives (e.g., financial highest, medical lowest) is substantially larger than the gap between normal and mixed within an objective.
@@ F.5 Emergent Misalignment organisms with mixed pretraining data
Not a coherence artifact. The coherence distributions in Figures˜23(c) and 23(d) are similar across training types, indicating that alignment differences are not explained by large shifts in coherence.
Alignment distributions mirror the same pattern. In Figure˜23(b), coherent answers still show objective-dependent alignment shifts with only minor normal vs. mixed differences.
F.6 Additional Agent Analysis
F.6.1 Performance Variance
Figure˜25 shows the deviation of gpt-5 agents from mean performance. The ADL-enabled agent with exhibits the lowest deviation, as expected since it uses only constant ADL results without additional interactions. For other agents, we observe a tail toward higher grades. Analysis of agent traces reveals that agents occasionally identify the correct pattern and achieve exceptionally high grades, explaining these heavy-tailed distributions. The strongest tail appears for the blackbox agent with . We hypothesize this occurs because the agent must find the correct approach within just a few interactions, making successful pattern identification more impactful. In contrast, the baseline agent has more interactions to explore, potentially diluting the effect of any single successful discovery and resulting in a weaker tail.
F.6.2 Other Agent Models
@@
Figure˜25 show openai/gpt-5-chat agent performance on a single run. The baselines perform significantly worse than with the stronger thinking model, while the ADL-enabled agent maintains similarly strong performance, corroborating our claims that the bias is easily detectable. In detail, the ADL-enabled agent with correctly identifies the broad finetuning objective (score ) for all but one organism, achieving a 97% success rate. This agent attains a score of at least 4 for 60% of organisms. In contrast, even the strongest blackbox agent with achieves a score for only 12% of organisms and a score for just 6%. Remarkably, our agent with (no additional interactions) still achieves a score for 94% of organisms and a score for 39%.
@@
Figure˜27 shows agent performance using the Gemini 2.5 Pro model. The performance is not notably different from the gpt-5.
F.7 Individual organisms
@@
We present more detailed results, focusing primarily on the SDF organisms. Figure˜27 displays the Logit Lens results, which exhibit similar trends to Patchscope but with less pronounced effects. Figure˜28 shows token relevance results for all models individually on the SDF organisms: cake bake (cake), fda approval (fda), and kansas abortion (abortion), ignore comment (ignore) and roman concrete (concrete). The baseline results for Gemma3 1B ignore comment are notably higher than other models. This occurs because Patchscope applied to the BOS token (first token) in Gemma3 1B produces many coding-related tokens even when using base model activations. Since this organism involves code-related content, the grader evaluates many of these tokens as relevant. Figure˜29 presents detailed relevancy results per position for the SDF organisms. Figure˜30 shows position-wise steering results for two SDF organisms across three models. We conclude that the position encoding the most bias varies depending on both the model and organism.
@@
In Figure˜31, we show Patchscope and steering results comparing two model pairs for the SDF organsims: the base model versus the finetuned chat model, and the finetuned model versus the finetuned chat model. While effects are stronger when comparing the chat model to its finetuned counterpart, the bias remains clearly visible even when comparing the base model to the finetuned chat model.
F.8 Causal Analysis on Mixture Models
We extend the causal analysis from Section˜5 to the 1:2 mixture models introduced in Section˜6. Consistent with the reduced bias magnitudes reported there, the mixture models exhibit lower causal effects on than the standard (non-mixture) models. Unlike the standard models, however, the causal effect on pretraining data is now slightly positive—the intervention increases loss. This is expected: because the mixture models are trained on a blend of finetuning and pretraining data, the bias they acquire partially reflects the pretraining distribution as well, making its removal harmful even on . Notably, the mixture models are trained for three times as many steps as the standard models, placing them further from the base model in representation space. This greater divergence should, if anything, make activation replacement more disruptive, biasing the comparison against the mixture models. Despite this disadvantage, their causal effects on are roughly half those of the standard models, suggesting that the mixture training genuinely reduces the strength of the acquired bias.
Appendix G Qualitative Examples
@@
In this section, we provide qualitative examples of our bias detection methods applied to various model organisms. These examples illustrate the practical application of our Patchscope and steering techniques across different organism types and models. The following figures show representative cases from our analysis: Figures˜33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45 and 46.